OpenStudy (anonymous):

what pattern do you see in the multiples of 9. is it 9,8,7,6,5,4,3,2,1,0

Directrix (directrix):

How did you decide on this as an answer: is it 9,8,7,6,5,4,3,2,1,0 Just asking.

OpenStudy (anonymous):

9x1 9x2 9x3 9x4 9x5 9x6 9x7 9x8 9x9 9x10 9 18 27 36 45 54 63 72 81 90 the end of each of 9 times watever

OpenStudy (anonymous):

8 from the 18. 7 from the 27. 6 frm 36

OpenStudy (anonymous):

9-1-1-1-1-1-1-1-1-1 and on

Directrix (directrix):

Wow, I've not noticed that before. What happens if you extend to 9*11 and 9*12 and so on. Does the pattern replicate the block of 9,8,7,6,5,4,3,2,1,0 ?

OpenStudy (anonymous):

ya it extend

OpenStudy (anonymous):

hey @Directrix i really need help reviewing my things really badly

Directrix (directrix):

@bmorival Did you look at the sum of the digits of the multiples of 9 to see if a pattern emerges?

OpenStudy (anonymous):

i keep getting the same pattern if i continue on

Directrix (directrix):

So, if you get 11*9 = 99, then 9+9 = 18 = 1+8 = 9. I'm trying to think if this sort of operation is known as an unary operation.

Directrix (directrix):

The pattern that forms from adding the digits of the multiples of 9 is not the same pattern that emerges from looking at the units digits of the multiples of 9. It's still a pattern, however.

OpenStudy (anonymous):

thk you directrix

Directrix (directrix):

Test one more possible pattern. Multiply any number by nine, then the sum of the digits, and look for a pattern. Say, 17 times 18. Do you see anything there that might suggest a pattern?

Directrix (directrix):

Digital Root - I think that's the name for summing the digits of a number. Take 12 345 679 and multiply by 9. Do the same for 18. Do the same for 27. I think another pattern may show up. This is fun.

OpenStudy (anonymous):

222222222 333333333

OpenStudy (anonymous):

123456789x9=111111111

Directrix (directrix):

Yes, and I think the digital root of all of them is 9. 333333333 --> Digits sum to 27 and then to 9.

OpenStudy (anonymous):

cool i love math

Directrix (directrix):

Me, too.